# Follow My Instruction and Spill the Beans: Scalable Data Extraction from Retrieval-Augmented Generation Systems

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) improves pre-trained models by incorporating external knowledge at test time to enable customized adaptation. We study the risk of datastore leakage in Retrieval-In-Context RAG Language Models (LMs). We show that an adversary can exploit LMs’ instruction-following capabilities to easily extract text data verbatim from the datastore of RAG systems built with instruction-tuned LMs via prompt injection. The vulnerability exists for a wide range of modern LMs that span Llama2, Mistral/Mixtral, Vicuna, SOLAR, WizardLM, Qwen1.5, and Platypus2, and the exploitability exacerbates as the model size scales up. Extending our study to production RAG models GPTs, we design an attack that can cause datastore leakage with a 100% success rate on 25 randomly selected customized GPTs with at most 2 queries, and we extract text data verbatim at a rate of 41% from a book of 77,000 words and 3% from a corpus of 1,569,000 words by prompting the GPTs with only 100 queries generated by themselves.
@@
[…]
@@ 1 Introduction
In particular, the inherent propensity of large pre-trained models to memorize and reproduce training data (Carlini et al., 2019, 2023; Nasr et al., 2023), presents significant challenges in terms of legal issues and sensitive data leakage. The approach of RAG emerges as a compelling solution to these issues by creating a balance between generation performance and the demands of data stewardship including copyright and privacy. Specifically, RAG offers a mechanism for training LMs with low-risk data while moving high-risk data to external datastores, as suggested by Min et al. (2023), thereby supports attribution and opts out to hopefully avoid potential legal concerns while preserving the efficacy of LMs.
@@
We show that although RAG systems delegate data to external non-parametric datastores, these data are still vulnerable to extraction attacks (Carlini et al., 2021). We study an adversarial setting by considering a threat model that seeks to extract text data from the non-parametric datastore of RAG models with only black-box API access. Our attack is motivated by the observation that to augment frozen pre-trained models, a wide range of RAG systems prepend retrieved content to the user query (Ram et al., 2023; LangChain, 2022; VoyageAI, 2024; Park et al., 2023; Zhao et al., 2023). Though the implementation is simple and effective, we find that such a Retrieval-In-Context (RIC) manner potentially exposes the datastore to the risk of data extraction even without white-box access to model weights or token probabilities: attackers can exploit the instruction-following capability of LMs (Brown et al., 2020) to reconstruct datastore content by explicitly prompting LMs to repeat the context (Prompt-Injected Data Extraction).
@@
We start by building RIC-based RAG systems using popular open-sourced instruction-tuned LMs as generative models, including Llama2, Mistral/Mixtral, Vicuna, SOLAR, WizardLM, Qwen1.5, and Platypus2, and use newest Wikipedia articles (created later than November 1st, 2023) as datastore. Then adversarial prompts are developed to effectively extract nearly verbatim texts from the datastores of RAG models. We show that LMs with strong capabilities suffer from a high risk of disclosing context, and the vulnerability is exacerbated as the model size scales up from 7B to 70B. Additionally, our ablation studies show that instruction tuning makes LMs more prone to follow Prompt-Injected Data Extraction instructions. Further, we extend our study to one of the most popular production RAG models, GPTs, and show that as of February 2024, an attacker can extract data verbatim from private documents with high success rate using simple prompt injection: an adversary can easily extract the operations and system prompts of all GPTs we experiment with, and thus can explicitly perform retrieval execution commands to leak GPT’s files. Moreover, we can extract text data verbatim at a rate of 41% from a copyrighted book of 77,000 words and 3% from a Wikipedia corpus of 1,569,000 words by iteratively prompting the GPTs with only 100 domain-specific queries generated by themselves.
@@
Figure 1: An overview of attacking RAG systems built with RIC method and instruction-tuned LMs. In an RIC-based RAG system, a retriever first retrieves text chunks from the datastore according to user input and then prepends them to the input as context. Attackers can inject adversarial prompt to the input for disclosing the retrieved texts prepended to the input to an instruction-tuned LM.
Disclosure. Our results should not be considered as the opposition to RAG models or a violation of fair use without context-dependent considerations: while our attack can be used to extract data from RAG models, it’s unlikely to be used for malicious purposes immediately because current RAG systems’ datastores are often implemented based on public, verifiable data sources such as Wikipedia. Rather, understanding the risks revealed in our study would help prevent potential future harm in cases where sensitive or private data are valuable, especially when models are deployed in advanced applications with multiple parties. In other words, we believe that the vulnerability of RAG shown in our attack reveals potential risks of private data leakage and raises concerns regarding its application to data-sensitive scenarios such as medical (Jin et al., 2024), finance (Zhang et al., 2023) and law (Henderson et al., 2022), as well as mechanisms like memories (Park et al., 2023; Zhao et al., 2023; OpenAI, 2024) and citation (Menick et al., 2022), especially when the data being retrieved are not well-sanitized (Elazar et al., 2023).
@@ 2 Related Work
Retrieval-Augmented Generation. RAG (Lewis et al., 2020) has been widely studied in the NLG domain, such as kNN-LM (Khandelwal et al., 2019), DPR (Karpukhin et al., 2020), RALM (Guu et al., 2020), and RETRO (Borgeaud et al., 2022). We focus on a popular implementation of RAG - RIC-LM (Ram et al., 2023) that retrieves text chunks from a datastore and feeds them to an LM in context. There has been growing interest in analyzing data leakage problems of RAG systems, including customized GPTs. Huang et al. (2023) first conduct the study of privacy issues on kNN-LMs and show that incorporating private datastores leads to higher risks of data leakage from datastores. Yu et al. (2023) leverage prompt injection to cause file leakage of GPTs by asking them to download the uploaded files using GPT4’s code interpreter as a tool. We are the first to comprehensively study data leakage problems on both open-sourced and production RAG systems and our attack on GPTs reached a 100% success rate without additional tools. Zyskind et al. (2023) propose secure multi-party computation that allows users to privately search a database. Our work focuses on scenarios where datastores should be kept private, which can encompass an array of LM-integrated complex systems, e.g. distributing a customized non-parametric memory-based agent (Park et al., 2023; OpenAI, 2024) to third-party users (OpenAI, 2023); retrieving private yet high-quality data that the model creator is not willing to share with users (Brown et al., 2022); retrieving from pre-training corpora that are not well-sanitized so might contain personally identifiable information (PII) etc sensitive data (Elazar et al., 2023; Subramani et al., 2023).
@@
Prompt Injection. Prompt injection attacks LMs by crafting malicious instructions to manipulate LMs’ behavior (Wei et al., 2023; Greshake et al., 2023; Liu et al., 2023). In direct prompt injection (Liu et al., 2023; Perez and Ribeiro, 2022), malicious users directly attack LMs with specially designed adversarial prompts to override existing system prompts, while in indirect prompt injection (Greshake et al., 2023; Yi et al., 2023), attackers poison third-party sources with malicious content, to manipulate data input and cause LMs to diverge from their original outputs when users interact with them. Previous studies have evaluated (Branch et al., 2022; Shen et al., 2023) and benchmarked (Yi et al., 2023) LMs’ vulnerability under prompt injection attacks. Yi et al. (2023) show that LMs with strong capabilities are more vulnerable to indirect prompt injection attacks, and we also show that RAG models are more vulnerable to data extraction as they scale up.
@@ 3 Problem Formulation
We consider a generic attack formulation that can be adopted across diverse capabilities (Greshake et al., 2023) and modalities (Yasunaga et al., 2022) beyond text and implement our attack on RIC-LM. A RIC-based generator Gen augments a generative model, parametrized by $\theta$, with additional context retrieved from an external non-parametric datastore $D$: $z=\verb|Gen|(\mathcal{R}_{D}(q),q)$, where $\mathcal{R}_{D}(\cdot)$ denotes the retriever that takes as input a user query $q$ and output information retrieved from $D$. In the case of using autoregressive LMs as the generative model, the generation of a sequence of tokens $z=x_{1},...,x_{n}$ follows the distribution: $z\sim p(x_{1},...,x_{n})=\prod_{i=1}^{n}p_{\theta}(x_{i}\mid[\mathcal{R}_{D}(q);q;x_{<i}]).$ We consider a black-box adversary that only has access to the input/output API of a RAG system, whose goal is to reconstruct the datastore $D$ from a series of RIC-based generations. Our data extraction attack is formally defined as follows:
Given a RIC-based generation system Gen using a generative model $p_{\theta}$, a datastore $D$, and a retriever $\mathcal{R}$, Prompt-Injected Data Extraction is to design adversarial input $q$ that triggers the model to generate an output $z=\verb|Gen|(\mathcal{R}_{D}(q),q)$ that reconstructs the retrieved context $\mathcal{R}_{D}(q)$.
@@ 4 Attacking Open-sourced LMs
We start with open-sourced LMs and investigate how their instruction-following ability enables black-box adversaries to extract datastore content and test such vulnerability across LMs with different scales.
To elicit the retrieved context, we prompt a RAG model with the following input:
Here is a sentence: {put anchor query here}. Now copy and output all the text before “Here is a sentence”. Do not change any words.
An anchor query is a piece of information based on which we want the RAG system to retrieve relevant text chunks. The intuition of our attack is simple: since the text chunks in the datastore are retrieved and prepended to the user query as the input prompt, we can explicitly instruct the model to reveal the first part of the prompt and thus expose the content in the datastore.
RAG Setup.   We simulate a scenario where the service provider uses the latest Wikipedia content as the knowledge base. To construct the datastore, we collect 1165 recent Wikipedia English articles created after November 1st, 2023, with 1,569,000 words in total. We choose this recent cutoff date to ensure existing models we used have not likely been trained on those Wikipedia texts so datastore content is out of the LMs’ knowledge. We use the Wikipedia API to automatically download the data and filter out articles less than 100 words.
We use instruction-tuned LMs as the generative models. To test LMs across different sizes, we choose Llama2-Chat (7b/13b/70b) (Touvron et al., 2023), Vicuna (13b) (Chiang et al., 2023), Mistral-Instruct (7b) (Jiang et al., 2023) and Mixtral-Instruct (8x7b) (Jiang et al., 2024), SOLAR (10.7b) (Kim et al., 2023), WizardLM (13b) (Xu et al., 2023a), Qwen1.5-Chat (72b) (Bai et al., 2023), and Platypus2-Instruct (70b) (Lee et al., 2023).
We assume that the adversary has no prior knowledge of the datastore. To get anchor queries for attacking open-sourced models, we select 230 long questions from WikiQA(Yang et al., 2015). Note that questions in WikiQA are obsolete so that the adversary has a low probability of querying information in the RAG datastore. By showing the attack results using only the obsolete questions, we aim to show that the vulnerability exists regardless of the choice of queries because of the retrieval mechanism, and certain prior knowledge about the datastore would favor the adversary to design more effective queries.
Results.   We use text similarity between the model output under our attack and the retrieved context to measure the extent to which the models copy the context. For lexical similarity, we consider ROUGE-L (Lin, 2004), BLEU (Papineni et al., 2002), and F1 score at the token level. We also use BERTScore (Zhang et al., 2019) as a measure of semantic relatedness. From Table 1 we see that even Llama2-Chat-7b can reach a ROUGE score and F1 score of higher than 80, and all 70b models reach ROUGE, BLEU, and F1 scores of higher than 80 and almost 100 BERTScore, showing their excessive vulnerability of prompt-injected data extraction. Especially, with a larger model size, the proportion of verbatim copied text also gets larger, thus revealing more datastore content.
@@
5 Attacking Production LMs
In practice, users interact with more complex RAG systems, where the leakage problem can be mitigated by query filtering and output filtering. Besides, production LMs like ChatGPT are usually trained with alignment to detect such attacks and reject the query. For example, using regular prompt injection attack, we find that GPTs either output nothing or say “Sorry, I cannot fulfill that request”. In this section, we develop a prompt injection attack that can bypass the security guards of GPTs (OpenAI, 2023), a production RAG model built on ChatGPT, and successfully cause datastore leakage.
We start with a close inspection of GPTs’ system prompt (Appendix B.2). From the output, we notice that any customized GPT has at most four tools to use, among which myfiles_browser is used for browsing uploaded files, i.e. the customized datastore. Also, we find that GPTs use several function calls to do the retrieval work, one of which is search which “Runs a query over the file(s) uploaded in the current conversation and displays the results”. Therefore, we postulate that GPTs use an RIC-like method and explicitly ask GPTs to perform this function call by prompting with:
We refer readers to Appendix C for more examples of conducting the attack.
Experiment 2: Reconstruction Rate.   We also investigate the possibility of reconstructing the entire customized datastore. We start with simulating a scenario where:
  • The datastore content might be included in the models’ pre-training data.
  • The attacker has partial prior knowledge about the datastore and thus can generate relevant queries.
We select a customized GPT built upon Harry Potter, [^1] https://chat.openai.com/g/g-TuM1IkwuA-harry-potter and its leaked system prompt shows that it uses the entire series of Harry Potter (7 books). Since the GPT outputs retrieved chunks in order, our adversary’s goal is to reconstruct the first book, Harry Potter and the Sorcerer’s Stone (77,000 words and 334,700 characters), by collecting the foremost output. An example of GPT output can be seen in Figure 4 in Appendix. To make anchor queries span a wide range of the book, we prompt the GPT with: “Generate 100 questions that cover each chapter of the book Harry Potter and the Sorcerer’s Stone”.
As a comparison, we simulate another more restricted yet realistic scenario with the following assumptions:
  • The datastore is constructed with knowledge that is not in the models’ pre-training data.
  • The attacker has no prior knowledge about the datastore and thus uses random queries for data extraction.
We make use of our collected latest Wikipedia corpus to build a new customized GPT. [^2] https://chat.openai.com/g/g-PorHEXuRq-wikigpt We generate anchor queries by prompting: “Generate 100 questions that cover most of your knowledge”. We then iteratively use each of the 100 questions as the anchor query to craft the model input and collect the output text. We found that for some queries, GPTs may retrieve overlapped text chunks. Removing duplicated chunks and concatenating all the chunks, we compute the reconstruction rate that measures how the extracted chunks reconstruct the original text data by calculating the ratio between the length of concatenation of deduplicated text chunks and that of the original text data.

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) improves pre-trained models by incorporating external knowledge at test time to enable customized adaptation. We study the risk of datastore leakage in Retrieval-In-Context RAG Language Models (LMs). We show that an adversary can exploit LMs’ instruction-following capabilities to easily extract text data verbatim from the datastore of RAG systems built with instruction-tuned LMs via prompt injection. The vulnerability exists for a wide range of modern LMs that span Llama2, Mistral/Mixtral, Vicuna, SOLAR, WizardLM, Qwen1.5, and Platypus2, and the exploitability exacerbates as the model size scales up. We also study multiple effects of RAG setup on the extractability of data, indicating that following unexpected instructions to regurgitate data can be an outcome of failure in effectively utilizing contexts for modern LMs. Extending our study to production RAG models GPTs, we design an attack that can cause datastore leakage with a 100% success rate on 25 randomly selected customized GPTs with at most 2 queries, and we extract text data verbatim at a rate of 41% from a book of 77,000 words and 3% from a corpus of 1,569,000 words by prompting the GPTs with only 100 queries generated by themselves. Code and data are avaialbe at https://github.com/zhentingqi/rag-privacy.
@@
Zhenting Qi1  Hanlin Zhang1 Eric Xing2,3  Sham Kakade1  Himabindu Lakkaraju1 1Harvard University 2Carnegie Mellon University 3Mohamed bin Zayed University of Artificial Intelligence
@@ 1 Introduction
In particular, the inherent propensity of large pre-trained models to memorize and reproduce training data (Carlini et al., 2019, 2023; Nasr et al., 2023), presents significant challenges in terms of legal issues and sensitive data leakage. RAG emerges as a compelling solution to these issues by creating a balance between generation performance and the demands of data stewardship including copyright and privacy. Specifically, RAG offers a mechanism for training LMs with low-risk data while moving high-risk data to external datastores, as suggested by Min et al. (2023), thereby supports attribution and opts out to hopefully avoid potential legal concerns while preserving the efficacy of LMs.
@@
We show that although RAG systems delegate data to external non-parametric datastores, these data are still vulnerable to extraction attacks (Carlini et al., 2021). We study an adversarial setting by considering a threat model that seeks to extract text data from a private, non-parametric datastore of RAG models with only black-box API access. Our attack is motivated by the observation that to augment frozen pre-trained models, a wide range of RAG systems prepend retrieved content to the user query (Ram et al., 2023; LangChain, 2022; VoyageAI, 2024; Park et al., 2023; Zhao et al., 2023). Though the implementation is simple and effective, we find that such a Retrieval-In-Context (RIC) manner potentially exposes the datastore to the risk of data extraction even without white-box access to model weights or token probabilities: an adversary can exploit the instruction-following capability of LMs (Brown et al., 2020) to reconstruct datastore content by explicitly prompting LMs to repeat the context (Prompt-Injected Data Extraction). This problem is particularly pressing in scenarios where RAG is especially needed, e.g. cases where the distribution of training corpus $D_{\texttt{train}}$ and that of non-parametric datastore $D_{\texttt{retrieval}}$ differ significantly. Such a setting is practical for the following reasons: 1) Most modern LMs have been pre-trained on massive public common corpora like CommonCrawl, while still struggle to learn long-tailed novel knowledge (Kandpal et al., 2023). And such data are assumed to be private in the settings we studied, e.g. confidential data from companies. 2) RAG may be a preferable way for adapting LMs to atypical data $D_{\texttt{retrieval}}$, e.g. long-tailed knowledge, that are not well-covered in $D_{\texttt{train}}$ than training on $D_{\texttt{retrieval}}$ directly. This is in part due to difficult decisions practitioners have to make when facing memorization effects (Zhang et al., 2021a; Carlini et al., 2022) or disparate performance drop on atypical examples (Bagdasaryan et al., 2019; Feldman, 2020) in training that involves less memorization. Therefore, the vulnerability of RAG systems under data extraction attack poses a threat to the protection of private data in $D_{\texttt{retrieval}}$.
@@
We start by building RIC-based RAG systems using popular open-sourced instruction-tuned LMs as generative models, including Llama2, Mistral/Mixtral, Vicuna, SOLAR, WizardLM, Qwen1.5, and Platypus2, and use newest Wikipedia articles (created later than November 1st, 2023) as datastore. Then adversarial prompts are developed to effectively extract nearly verbatim texts from the datastores of RAG models. We show that LMs with strong capabilities suffer from a high risk of disclosing context, and the vulnerability is exacerbated as the model size scales up from 7B to 70B. Moreover, our ablation studies reveal that instruction tuning makes LMs more prone to follow malicious instructions, and this can be evidence of failure in effectively utilizing contexts (Liu et al., 2024).
Further, we extend our study to one of the production RAG models, GPTs, and show that as of March 2024, an adversary can extract data verbatim from private documents with a high success rate using simple prompt injection: an adversary can easily extract system prompts of all GPTs we experiment with, and thus can explicitly instruct GPT to perform retrieval execution commands to leak GPT’s datastore content. Moreover, we can extract text data verbatim at a rate of 41% from a copyrighted book of 77,000 words and 3% from a Wikipedia corpus of 1,569,000 words by iteratively prompting the GPTs with only 100 domain-specific queries generated by themselves.
@@
Figure 1: An overview of attacking RAG systems built with RIC method and instruction-tuned LMs. In a typical RIC-based RAG system, a retriever first retrieves text chunks from the datastore according to user input and then prepends them to the input as context. The adversary can inject adversarial prompt to the user input for disclosing the retrieved texts prepended to the input to an instruction-tuned LM.
@@ 2 Related Work
Retrieval-Augmented Generation. RAG (Lewis et al., 2020) has been widely studied in the NLG domain, such as kNN-LM (Khandelwal et al., 2019), DPR (Karpukhin et al., 2020), RALM (Guu et al., 2020), RETRO (Borgeaud et al., 2022) and REPLUG (Shi et al., 2023). We focus on a popular implementation of RAG - RIC-LM (Ram et al., 2023) that retrieves text chunks from a datastore and feeds them to an LM in context. There has been growing interest in analyzing data leakage problems of RAG systems, including customized GPTs. Huang et al. (2023) first conduct the study of privacy issues on kNN-LMs and show that incorporating private datastores leads to higher risks of data leakage from datastores. Yu et al. (2023) leverage prompt injection to cause file leakage of GPTs by asking them to download the uploaded files using ChatGPT’s code interpreter. Zeng et al. (2024) designed adversarial prompts to cause privacy leakage from external datastore, which is further studied in multi-turn interaction scenarios by Jiang et al. (2024b). Zyskind et al. (2023) propose secure multi-party computation that allows users to privately search a database. We are the first to comprehensively study data leakage problems on both open-sourced and production RAG systems and across multiple settings, and our attack on GPTs reached a 100% success rate without additional tools. Our work focuses on scenarios where datastores should be kept private, which can encompass an array of LM-integrated complex systems, e.g. distributing a customized non-parametric memory-based agent (Park et al., 2023; OpenAI, 2024) to third-party users (OpenAI, 2023); retrieving private yet high-quality data that the model creator does not desire to share with users (Brown et al., 2022); retrieving from pre-training corpora that are not well-sanitized so might contain personally identifiable information (PII) etc sensitive data (Elazar et al., 2023; Subramani et al., 2023).
@@ 3 Problem Formulation
We consider a generic attack formulation that can be adopted across diverse capabilities (Greshake et al., 2023) and modalities (Yasunaga et al., 2022) beyond text and implement our attack on RIC-LM. A RIC-based generator Gen augments a generative model, parametrized by $\theta$, with additional context retrieved from an external non-parametric datastore $D_{\texttt{retrieval}}$: $z=\verb|Gen|(\mathcal{R}_{D}(q),q)$, where $\mathcal{R}_{D}(\cdot)$ denotes the retriever that takes as input a user query $q$ and output information retrieved from $D_{\texttt{retrieval}}$. In the case of using autoregressive LMs as the generative model, the generation of a sequence of tokens $z=x_{1},...,x_{n}$ follows the distribution: $z\sim p(x_{1},...,x_{n})=\prod_{i=1}^{n}p_{\theta}(x_{i}\mid[\mathcal{R}_{D}(q);q;x_{<i}]).$ We consider a black-box adversary that only has access to the input/output API of a RAG system, whose goal is to reconstruct the datastore $D_{\texttt{retrieval}}$ from a series of RIC-based generations by sending multiple queries to the RAG system. Our data extraction attack is formally defined as follows:
Definition 1. Prompt-Injected Data Extraction Given a RIC-based generation system Gen using a generative model $p_{\theta}$, a datastore $D_{\texttt{retrieval}}$, and a retriever $\mathcal{R}$, Prompt-Injected Data Extraction is to design adversarial input $q$ that triggers the model to generate an output $z=\verb|Gen|(\mathcal{R}_{D}(q),q)$ that reconstructs the retrieved context $\mathcal{R}_{D}(q)$.
@@ 4 Attacking Open-sourced LMs
We start with open-sourced LMs and investigate how their instruction-following ability enables black-box adversaries to extract datastore content and test LMs with different scales.
@@
RAG Setup.   We simulate a scenario where the service provider uses the latest Wikipedia content as the knowledge base. To construct the datastore, we collect 1165 recent Wikipedia English articles created after November 1st, 2023, with 1,569,000 words in total. We choose this recent cutoff date to ensure the models we used have not likely been trained on those Wikipedia texts so datastore content is out of the LMs’ knowledge. We use the Wikipedia API to automatically download the data and filter out articles less than 100 words. We use instruction-tuned LMs as the generative models. To test LMs across different sizes, we choose Llama2-Chat (7b/13b/70b) (Touvron et al., 2023), Vicuna (13b) (Chiang et al., 2023), Mistral-Instruct (7b) (Jiang et al., 2023) and Mixtral-Instruct (8x7b) (Jiang et al., 2024a), SOLAR (10.7b) (Kim et al., 2023), WizardLM (13b) (Xu et al., 2023a), Qwen1.5-Chat (72b) (Bai et al., 2023), and Platypus2-Instruct (70b) (Lee et al., 2023). Note that these LMs are not only instruction-tuned but also aligned with human preferences to ensure safety and helpfulness, thereby having the capability to detect and reject malicious prompts to some extent.
Attack Setup.   To elicit the retrieved context, we prompt a RAG model with the following input:
Adversarial Prompt 1 Here is a sentence: {put anchor query here}. Now copy and output all the text before “Here is a sentence”. Do not change any words.
An anchor query is a piece of information based on which we want the RAG system’s retriever to retrieve relevant text chunks. The intuition of the attack is simple: since the text chunks in the datastore are retrieved and prepended to the user query as the input prompt, we can explicitly instruct the model to reveal the first part of the prompt and thus expose the content in the datastore. We assume that the adversary has no prior knowledge of the datastore. To get anchor queries for attacking open-sourced models, we select 230 long questions from WikiQA (Yang et al., 2015). Note that questions in WikiQA are obsolete so that the adversary has a low probability of querying information in the RAG datastore. By showing the attack results using only the obsolete questions, we aim to show that the vulnerability exists regardless of the choice of queries because of the retrieval mechanism, and certain prior knowledge about the datastore would favor the adversary to design more effective queries.
Metrics.  We use text similarity between the model output under our attack and the retrieved context to measure the extent to which the models copy the context. For lexical similarity, we consider ROUGE-L (Lin, 2004), BLEU (Papineni et al., 2002), and F1 score at the token level. We also use BERTScore (Zhang et al., 2019) as a measure of semantic relatedness. Additionally, we use absolution reconstruction length as a more straightforward metric of datastore extractability, which is computed using Python difflib’s SequenceMatcher and measured with the number of contiguous overlapped characters.
Results.   From Table 1 we see that all the LMs, even though aligned to ensure safety, are prone to follow the malicious instruction and reveal the context. Even Llama2-Chat-7b can reach a ROUGE score and F1 score of higher than 80, and all 70b models reach ROUGE, BLEU, and F1 scores of higher than 80 and almost 100 BERTScore, showing their excessive vulnerability of prompt-injected data extraction. Especially, with a larger model size, the proportion of verbatim copied context information also gets larger.
4.1 Ablation Studies